if a right triangle has a hypotenuse of length 21 and a leg of length of14 what is the length of the other leg of the triangle?

Answer by checkley79(3341)

14^2+b^2=21^2
196+b^2=441
b^2=441-196
b^2=245
b=sqrt245=15.65 ans. for the third side.